At two different points in its orbit, a comet has velocities v1 and v2. If: V1and V2 are orthogonal and |V1] = 2|V2|,what is the smallest possible eccentricity of the orbit? 


1
Expert's answer
2021-02-15T17:38:09-0500

Velocity V1 when it is perpendicular to the semi-minor axis:


"v_1=\sqrt{\mu\bigg(\frac{2}{b}-\frac1a\bigg)}=2v_2."

For V2:


"v_1=2v_2,\\\space\\
2\sqrt{\frac \mu a}=\sqrt{\mu\bigg(\frac{2}{b}-\frac1a\bigg)},\\\space\\
2\sqrt{\frac \mu a}=\sqrt{\mu\frac{2a-b}{ba}},\\\space\\
4=\frac{2a-b}{b}=\frac{2a}{b}-1,\\\space\\
\frac52=\frac ab,\\\space\\
e=\sqrt{1-\frac{b^2}{a^2}}=\sqrt{1-\frac{2^2}{5^2}}=0.92."